OpenStudy (anonymous):

Sam has a total of 58 DVDs and CDs. If the number of CDs is two more than three times the number of DVDs how many CDs does he have?

OpenStudy (anonymous):

no of cds:44 no of dvds:14

OpenStudy (anonymous):

Let the number of DVDs be d and the number of CDs be c. c+d=58 c=2+3d Solve for c and d.

OpenStudy (anonymous):

and how do you do that?

OpenStudy (anonymous):

Substitute c=2+3d. Then you can work out d. And then substitute d into c=2+3d.

OpenStudy (anonymous):

I still don't understand

OpenStudy (anonymous):

I know your supposed to move something around but what?

OpenStudy (anonymous):

Now you have two equations: c+d=58 c=2+3d If you substitute c=2+3d into c+d=58, then the first equation will be like: 2+3d+d=58, which is 4d=58-2=56 So d=56/4=14